how do i find the SLOPE of
a. (8,3), (-4,3)
b. (9,-2), (3,4)

Answer by fractalier(6550)

Use the slope formula
m = (y2 - y1) / (x2 - x1)
for the first pair we have
m = (3 - 3) / (-4 - 8) = 0
and for the second
m = (4 - -2) / (3 - 9) = 6 / -6 = -1